OpenStudy (anonymous):

in a game a coin is flipped twice, if heads lands up either time, the player wins $4.60 otherwise the player looses $6.00. What is the expected value of this game?

OpenStudy (anonymous):

4 equally likely situations here HH: win 4.60 HT: win 4.60 TH: win 4.60 TT: lose 6 Expected value = 3/4*(4.60) + 1/4*(-6)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you can also do this as follow: play the game 4 times, you win \(3\times 4.60\) and lost \(-6\) for a total of \(3\times \$4.60-\$6\) averaged over the 4 games, it is \[\frac{3+4.60-6}{4}\]